OpenStudy (anonymous):

factor za^2 - 2a - 2z^a + 4z

OpenStudy (anonymous):

factor out za from the first and third terms, factor out 2 from second and fourth, and there should be something more that you can factor :P

OpenStudy (anonymous):

In the third term, z is raised to a. How can I factor out za from the third term? Thanks for the help. :)

OpenStudy (amistre64):

hmmm wondering if we can do a substitution

OpenStudy (amistre64):

if it was factorable, we could equate it to 0 za^2 - 2a - 2z^a + 4z = 0 a(za - 2) - 2z (z^(a-1) -2) = 0 a(za - 2) = 2z (z^(a-1) -2) a(za - 2)/z = 2(z^(a-1) -2) not real sure that it can be facotred

OpenStudy (amistre64):

completing the square in "a" maybe

OpenStudy (anonymous):

Thanks. :)

OpenStudy (amistre64):

\[za^2 - 2a - 2z^a + 4z \] \[z(a^2 - 2a +1) -z- 2z^a + 4z \] \[z(a-1)^2 - 2z^a +3z \] assuming that z not equal to 0 \[z(a-1)^2 - 2z^a +3z =0\] \[(a-1)^2 - 2z^{a-1} +3 =0\] \[(a-1)^2= 2z^{a-1} -3\] \[a-1= \pm~\sqrt{2z^{a-1} -3}\] \[a= 1\pm~\sqrt{2z^{a-1} -3}\] but thats not factoring it; its just trying to define the roots